2. In a poker hand consisting of 5 cards, find the probability of holding
a. 3 kings
b. 4 hearts and 1 diamonds
1
Expert's answer
2020-10-19T17:46:54-0400

A poker hand consists of 5 cards. The sample space has "\dbinom{52}{5}" possible outcomes.

a. Choose 3 kings from 4 kings. From the remaining "52-4=48" cards choose 2 non-kings to make a five-card poker hand

"P(3\ kings)=\dfrac{\dbinom{4}{3}\dbinom{52-4}{5-3}}{\dbinom{52}{5}}="

"=\dfrac{\dfrac{4!}{3!(4-3)!}\cdot\dfrac{48!}{2!(48-2)!}}{\dfrac{52!}{5!(52-5)!}}="

"=\dfrac{2(48)(47)(120)}{52(51)(50)(49)(48)}=\dfrac{94}{54145}\approx0.001736"

b. There are total 13 hearts and 13 diamonds.

Choose 4 hearts from 13 hearts


"\dbinom{13}{4}"

Choose 1 diamonds from 13 diamonds


"\dbinom{13}{1}"


"P(4\ hearts\ and\ 1\ diamonds)=\dfrac{\dbinom{13}{4}\dbinom{13}{1}}{\dbinom{52}{5}}="

"=\dfrac{\dfrac{13!}{4!(13-4)!}\cdot\dfrac{13!}{1!(13-1)!}}{\dfrac{52!}{5!(52-5)!}}="

"=\dfrac{13(12)(11)(10)(5)(13)}{52(51)(50)(49)(48)}=\dfrac{143}{39984}\approx0.003576"